8 workers can paint a building in 24 days.How many days will 18 workers take to paint the same building
Answer by Theo(13342)

this can be an inverse variation type problem.
the formula to use is y = k/x
set y = 24 days and set x = 8 workers and solve for k.
you get 24 = k/8
solve for k to get k = 192

k is the constant of variation and will always remain the same.
when x = 18, the formula becomes y = 192/18.
solve for y to get y = 10 and 2/3 days.

to see if this is correct, use the rate of each person * number of people * time = quantity of work formula.

when people paint 1 building in 24 days, this formula becomes:

r * 8 * 24 = 1

r is the rate of each person.
8 is the number of people.
24 is the number of days.
1 is the quantity of work produced (1 painted building).

solve for r to get r = 1 / (8 * 24)
you get r = 1/192.

when the number of people is 18 and the rate of each person is 1/192, the formula becomes:

1/192 * 18 * t = 1

1/192 is the rate of each person.
this means each person paints 1/192 of the building in one day.
18 is the number of people painting.
t is the number of days
1 is the quantity of work produced (1 painted building).

solve for t to get t = 1 / (18 * 1/192)
simplify to get t = 1 / (18/192)
this is equivalent to t = 1 * 192/18
simplify to get t = 192/18
convert to a mixed fraction to get t = 10 and 2/3 days.

the answer is the same whether you use the inverse variation formula or the rate of each person * number of people * time = quantity of work produced formula.

your solution should be 10 and 2/3 days.